# Open Packing in Graphs: Bounds and Complexity

### v1

@@ 3.2 Subclasses of $(P_{4}\cup rK_{1})$-free Graphs
Let $H$ be a graph on $p$ vertices for some $p\geq 4$ such that Open Packing is polynomial time solvable on $H$-free graphs. Then, we prove that $H\in\{P_{4}\cup(p-4)K_{1},P_{3}\cup(p-3)K_{1},K_{2}\cup(p-2)K_{1},pK_{1}\}$ unless $P=NP$. On the contrary, assume that $H\notin\{P_{4}\cup(p-4)K_{1},P_{3}\cup(p-3)K_{1},K_{2}\cup(p-2)K_{1},pK_{1}\}$. Then, by Observation 1, $H$ contains at least one of (i) $K_{3}$ (ii) $2K_{2}$ (iii) $C_{4}$ (iv) $K_{1,3}$ or (v) $C_{5}$ as its induced subgraph. So, (a) $H^{\prime}$-free graphs are subclass of $H$-free graphs for some $H^{\prime}\in\{K_{3},2K_{2},C_{4},K_{1,3},C_{5}\}$ and (b) Open Packing is NP-complete on $H^{\prime}$-free graphs for every $H^{\prime}\in\{K_{3},2K_{2},C_{4},K_{1,3},C_{5}\}$ (by Remark 1 and Theorem 3). Thus by (a), (b) and the fact that Open Packing is in NP in every subclass of simple graphs, we can conclude that Open Packing is NP-complete on $H$-free graphs, a contradiction. This proves the necessary of Theorem 2.
The sufficiency part of Theorem 2 holds by Theorem 8. ∎
Note that Observation 1 along with Theorem 7 and that Total Dominating Set is NP-complete for (i) Split graphs ($\{2K_{2},C_{4},C_{5}\}$-free graphs) [1], (ii) Bipartite graphs (a subclass of $K_{3}$-free graphs) [16] and (iii) Line graphs (a subclass of $K_{1,3}$-free graphs) [14] implies the following theorem similar to that of Theorem 2 for Total Dominating Set.
@@
For $p\geq 4$, let $H$ be a graph on $p$ vertices. Then, Total Dominating Set is polynomial time solvable on the class of $H$-free graphs if and only if $H\in\{pK_{1},(K_{2}\cup(p-2)K_{1}),(P_{3}\cup(p-3)K_{1}),(P_{4}\cup(p-4)K_{1})\}$ unless $NP=P$.
@@ 3.3 $H$ on three vertices
Note that $3K_{1},(K_{2}\cup K_{1}),P_{3}$ and $K_{3}$ are the only graph on three vertices. By Theorem 8, an optimal open packing of a $H$-free graph can be found in polynomial time if $H\in\{3K_{1},(K_{2}\cup K_{1}),P_{3}\}$ and Open Packing is NP-complete on $K_{3}$-free graphs by Remark 1. This completes the proof of Theorem 1. ∎
Note that (a) Total Dominating Set is NP-complete for $K_{3}$-free graphs [16] and (b) Theorem 7 proves a result similar to that of Theorem 1 for Total Dominating Set. This note completes our study on the complexity (P vs NPC) of Total Dominating Set and Open Packing on $H$-free graphs. Next, we study the complexity of these problems on some subclasses of split graphs.
@@ 4.1 $K_{1,r}$-free Split Graphs
We begin this section with a construction to prove that Open Packing is NP-complete in $K_{1,4}$-free split graphs.
@@
Step 1 : Subdivide each edge $e=uu^{\prime}$ in $G$ into a three vertex path $ueu^{\prime}$ in $G^{\prime}$. Step 2 : Introduce three new vertices $x$, $y$, $z$ and two edges $xy,xz$ in $G^{\prime}$. Step 3 : Make $E(G)\cup\{y,z\}$ a clique in $G^{\prime}$.
@@
On the contrary, assume that $\{w,w_{1},w_{2},w_{3},w_{4}\}$ induces a $K_{1,4}$ in $G^{\prime}$ with $w$ as its centre. Then, $\{w_{1},w_{2},w_{3},w_{4}\}$ is an independent set in $G^{\prime}$. So $w\notin I$, since $N_{G^{\prime}}(v)\subseteq C$ is a clique for every vertex $v\in I$ whereas $N_{G^{\prime}}(w)$ contains an independent set on four vertices. This implies that $w\in C$. Note that by Construction 2, for every vertex $v\in C$, $|N_{G^{\prime}}(v)\cap I|\leq 2$ and so $w$ has at most two neighbours in $I$. Therefore, |$|\{w_{1},w_{2},w_{3},w_{4}\}\cap C|\geq 2$, a contradiction to $C$ being a clique and $\{w_{1},w_{2},w_{3},w_{4}\}$ is an independent set in $G^{\prime}$. Thus, $G^{\prime}$ is a $K_{1,4}$-free graph.
The following claims are useful to prove the guarantee of Construction 2.
@@
Proof of Claim 7 is similar to the Proof of Claim 1 and so it is moved to Appendix C
@@
Observe that in $G^{\prime}$, the graph induced by $C\cup\{x\}$ consists of a complete graph ($G^{\prime}[C]$) on $m+2$ vertices together with a vertex $x$ which is adjacent to exactly two vertices ($y\text{ and }z$) in $C$. So any two distinct vertices in $C\cup\{x\}$ have a common neighbour. Thus, the claim holds.
@@
The following theorem holds by Construction 2 and the fact that Independent Set is NP-complete on simple graphs [12].
@@
Similar to that of Theorems 5 and 6, it can be proved that on $K_{1,4}$-free split graphs (i) Open Packing parameterized by solution size is W[1]-complete and (ii) Max-Open Packing is hard to approximate within a factor of $N^{(\frac{1}{2}-\epsilon)}$ for any $\epsilon>0$ unless $NP=P$, where $N$ denotes the number of vertices in a $K_{1,4}$-free split graph (proof of (ii) is given as Theorem 16 in Appendix C).
Next, to complete the dichotomy result of Open Packing in $K_{1,r}$-free split graphs, we prove that a maximum open packing in $K_{1,3}$-free split graphs can be found in polynomial time and the following claim and lemma are useful in proving it.
@@
Let $S$ be an open packing in a connected split graph $G(C\cup I,E)$ with $S\cap C\neq\emptyset$. Let $u\in S\cap C$. Then, we show that $S\subseteq N[u]$. On the contrary, assume that there exists a vertex $v\in S\setminus N[u]$. Since $C\subseteq N[u]$ and $v\notin N[u]$, $v\notin C$, i.e., $v\in I$. Since $G$ is connected, $deg(v)\geq 1$ and so there exists $w\in C$ such that $vw\in E(G)$. Since $v\notin N(u)$, $w\neq u$. Further since $w,u\in C$, $w\in N(u)$. This implies that $w\in N(u)\cap N(v)$, a contradiction to $S$ is an open packing in $G$ and $u,v\in S$. Therefore, $S\subseteq N[u]$. Since $S$ is an open packing in $G$, $|S\cap N(u)|\leq 1$. So, $|S|\leq|S\cap N(u)|+1\leq 2$.
@@
Let $I=\{x_{1},x_{2},\ldots,x_{k}\}$. Let $V_{i}=N(x_{i})\cap C$ for $i=1,2\ldots,k$ and $V_{0}=C\setminus(\cup_{i=1}^{k}V_{i})$. Then, $V_{i}\cap V_{j}=\emptyset$ for every $i,j\in\{1,2,\ldots,k\}$ and $i\neq j$. If not, let $u^{\prime}\in V_{i}\cap V_{j}$ for some $i,j\in\{1,2,\ldots,k\}$ with $i\neq j$. Then, $x_{i},x_{j}\in N(u^{\prime})\cap I$, a contradiction to $|N(u)\cap I|\leq 1$ for every $u\in C$. We prove that in this case $\rho^{o}(G)\in\{2,|I|\}$ in two parts.
@@
Note that the assumption on Case 1.2 is the negation of Case 1.1. As we have already proved that $N(x_{i})\cap N(x_{j})=V_{i}\cap V_{j}=\emptyset$ for every $x_{i},x_{j}\in I$, $I$ is an open packing in $G$, and hence $\rho^{o}(G)\geq|I|=k$. Next, we show that $\rho^{o}(G)\leq|I|=k$. On the contrary, assume that $\rho^{o}(G)>k$. Let $S$ be an open packing in $G$ of size $\rho^{o}(G)$. Then, $|S|>|I|=k$, and so $S\cap C\neq\emptyset$. Therefore, by Claim 9, $|S|\leq 2$. This implies that $2\geq|S|>|I|\geq 1$. So, $|S|=2$ and $|I|=k=1$. Then, $V(G)=C\cup\{x_{1}\}$ and by assumption on Case 1.2, $deg(x_{1})\geq 2$. Let $S=\{a,b\}$. Since, $G$ is constituted of a complete graph ($G[C]$) together with a vertex $x_{1}$ which is adjacent to at least two vertices in $C$, for any choice of $a,b\in C\cup\{x_{1}\}$, $a$ and $b$ will have a common neighbour in $G$, a contradiction to $S$ is an open packing of size two in $G$. Therefore, for every open packing $S$ in $G$, $|S|\leq|I|=k$. This implies that $\rho^{o}(G)\leq|I|$. Since we have proved that $I$ is an open packing in $G$, $\rho^{o}(G)=|I|$.
@@
If there exist $x,y\in I$ such that $N(x)\cap N(y)=\emptyset$, then $\{x,y\}$ is an open packing of size two in $G$, and so $\rho^{o}(G)=2$. Else, there exists a vertex $x\in I$ such that $deg(x)=1$. Let $N(x)=\{w\}$. Then, it is evident that $\{x,w\}$ is an open packing of size two. Thus, $\rho^{o}(G)=2$.
@@
Note that the assumption on Case 2.2 is the negation of Case 2.1. We prove that $\rho^{o}(G)=1$ in this case. As we have already proved that $\rho^{o}(G)\leq 2$, it is enough to rule out the case $\rho^{o}(G)=2$. Since $V(G)\subseteq C\cup I$, where (i) any two vertices in $I$ have a common neighbour (ii) every vertex $x\in I$ has at least two neighbours in $C$ and (iii) $G[C]$ is a complete graph, any two vertices in $V(G)$ have a common neighbour in $G$ and so $\rho^{o}(G)<2$. Therefore, $\rho^{o}(G)=1$.
Hence, we have shown every possible value of $\rho^{o}(G)$. Note that every $K_{1,3}$-free split graph $G$ falls under Case 1 or Case 2. It can be easily verified in polynomial time whether a $K_{1,3}$-free split graph $G$ falls under Case 1 or Case 2. Also, the conditions given in Cases 1.1, 1.2, 2.1 and 2.2 are polynomial time verifiable. Hence, the open packing number of a $K_{1,3}$-free split graph can be found in polynomial time. ∎
@@ 4.2 $I_{r}$-split graphs
We observed that a minor modification in Corneil and Perl’s reduction [1] will show that Total Dominating Set is NP-complete on $I_{2}$-split graphs. We extended this by proving a dichotomy that Total Dominating Set in $I_{r}$-split graphs is (i) NP-complete for $r\geq 2$ and (ii) polynomial time solvable for $r=1$. In addition, we prove that Open Packing is (i) NP-complete on $I_{r}$-split graphs for $r\geq 3$ and (ii) polynomial time solvable on $I_{r}$-split graphs for $r\leq 2$. We begin the section with Construction 3 to prove that Total Dominating Set is NP-complete on $I_{r}$-split graphs for $r\geq 2$.
@@
Figure 8: (a) An instance $(U,\mathcal{W})$ of 3-Hitting Set problem and (b) a $I_{3}$-split graph $G(C\cup I,E)$ produced from Construction 3 where $C=U\cup\{x_{1},x_{2},x_{3}\}$ is a clique in $G$ and $I=\{z_{W_{1}},z_{W_{2}},z_{W_{3}},z_{W_{4}},y\}$ is an independent set in $G$ with $deg_{G}(z_{W_{i}})=deg_{G}(y)=3$ for every $i=1,2,3,4$ and $V(G)=C\cup I$.
An example of Construction 3 for the case $r=3$ is given in Fig. 8. Note that $V(G)=C\cup I$, where $C=U\cup\{x_{1},x_{2},\ldots,x_{r}\}$ and $I=\{z_{W}\,:\,W\in\mathcal{W}\}\cup\{y\}$ and $E(G)=\{uv:u,v\in U\cup\{x_{1},x_{2},\ldots,x_{r}\}\}\cup\{uz_{W}\,:\,u\in U,W\in\mathcal{W}\text{ and }u\in W\}\cup\{x_{i}y\,:\,1\leq i\leq r\}$. So, $|V(G)|=|U|+|W|+r+1$ and $|E(G)|=\binom{|U|+r}{2}+(r\cdot|W|)+r$. Hence, the graph $G$ can be constructed in polynomial time with respect to $|U|,|\mathcal{W}|$ and $r$. By the construction of $G$, it is clear that $V(G)=C\cup I$ is a clique-independent set partition of the vertex set of $G$. So, $G$ is a split graph. Since the size of every set $W\in\mathcal{W}$ is $r$, $deg_{G}(z_{W})=r$ by the construction of $G$. Further, since $N_{G}(y)=\{x_{1},x_{2},\ldots,x_{r}\}$, $deg_{G}(y)=r$. This implies that $deg_{G}(u)=r$ for every $u\in I$. Thus, $G(C\cup I,E)$ is an $I_{r}$-split graph.
@@
Let $D$ be a total dominating set of size $k+1$ in $G$. Then, $D\cap\{x_{1},x_{2},\ldots,x_{r}\}\neq\emptyset$ as $N_{G}(y)=\{x_{1},x_{2},\ldots,x_{r}\}$ and $D\cap N_{G}(y)\neq\emptyset$. Let $D^{\prime}=D\cap U$. Since $D\cap\{x_{1},x_{2},\ldots,x_{r}\}\neq\emptyset$ and $U\cap\{x_{1},x_{2},\ldots,x_{r}\}=\emptyset$, $|D^{\prime}|=|D\cap U|\leq k$. Next, we show that $D^{\prime}$ is a hitting set of size at most $k$ of $(U,\mathcal{W})$. Let $W\in\mathcal{W}$. Then, since $D$ is a total dominating set, there exists a vertex $u\in D$ such that $uz_{W}\in E(G)$. Also, since $N_{G}(z_{W})\subseteq U$ by the construction of $G$, $u\in D\cap U=D^{\prime}$. So, by the construction of $G$, $u\in W$. Since $W$ is an arbitrary set in $\mathcal{W}$, for every $W\in\mathcal{W}$, there exists an element $u\in D^{\prime}$ such that $u\in W$. Hence, $D^{\prime}$ is a hitting set of size at most $k$ in $(U,\mathcal{W})$. ∎
@@
The following theorem completes the dichotomy result for Total Dominating Set in $I_{r}$-split graphs.
@@
Let $N(I)=\{u\}$. Since $I$ is an independent set in $G$, $u\in C$. Since $C$ is a clique and $N(I)=\{u\}$, every vertex in $V(G)\setminus\{u\}=(C\cup I)\setminus\{u\}$ is adjacent to $u$. Let $v\in I$. Then, since $N(I)=\{u\}$, $vu\in E(G)$. This implies that every vertex in $V(G)=(V(G)\setminus\{u\})\cup\{u\}$ is adjacent to a vertex in $\{u,v\}$. So, $\{u,v\}$ is a total dominating set in $G$ and hence $\gamma_{t}(G)\leq 2$. Since $\gamma_{t}(H)\geq 2$ for every graph $H$ that admits a total dominating set, $\gamma_{t}(G)=2$.
@@
Since $I$ is an independent set in $G$, $N(I)\subseteq C$. Let $u,u^{\prime}\in N(I)$. Then, since $C$ is a clique and $u,u^{\prime}\in N(I)\subseteq C$, every vertex in $C$ is adjacent to $u$ or $u^{\prime}$. Note that by the definition of $N(I)$, every vertex in $I$ is adjacent to some vertex in $N(I)$. Thus, every vertex in $V(G)=C\cup I$ is adjacent to a vertex in $N(I)$ and so $N(I)$ is a total dominating set. This implies that $\gamma_{t}(G)\leq|N(I)|$. But since $\gamma_{t}(G)\geq|N(I)|$, we conclude that $\gamma_{t}(G)=|N(I)|$.
Thus by Cases 1 and 2, $\gamma_{t}(G)=\max\{2,|N(I)|\}$. Note that $|N(I)|$ can be found in linear time. So, an optimal total dominating set of $G$ can be found in linear time. ∎
Next, we prove the dichotomy result on Open Packing in $I_{r}$-split graphs. Firstly, we show that Open Packing is NP-complete on $I_{r}$-split graphs for $r\geq 3$.
@@
Input: A collection of sets $X_{1},X_{2},\ldots,X_{r}$ such that $|X_{i}|=q$ for $i=1,2\ldots,r$, for some $q\in\mathbb{N}$ and a non-empty set $M\subseteq\prod\limits_{i=1}^{r}X_{i}$.

### v2

@@ 3.2 Subclasses of $(P_{4}\cup rK_{1})$-free Graphs
Let $H$ be a graph on $p$ vertices for some $p\geq 4$ such that Open Packing is polynomial time solvable on $H$-free graphs. Then, we prove that $H\in\{P_{4}\cup(p-4)K_{1},P_{3}\cup(p-3)K_{1},K_{2}\cup(p-2)K_{1},pK_{1}\}$ unless P = NP. On the contrary, assume that $H\notin\{P_{4}\cup(p-4)K_{1},P_{3}\cup(p-3)K_{1},K_{2}\cup(p-2)K_{1},pK_{1}\}$. Then, by Observation 1, $H$ contains at least one of (i) $K_{3}$, (ii) $2K_{2}$, (iii) $C_{4}$, (iv) $K_{1,3}$, or (v) $C_{5}$ as its induced subgraph. So, (a) $H^{\prime}$-free graphs are a subclass of $H$-free graphs for some $H^{\prime}\in\{K_{3},2K_{2},C_{4},K_{1,3},C_{5}\}$ and (b) Open Packing is NP-complete on $H^{\prime}$-free graphs for every $H^{\prime}\in\{K_{3},2K_{2},C_{4},K_{1,3},C_{5}\}$ (by Remark 1 and Theorem 3). Thus, by (a), (b), and the fact that Open Packing is in NP in every subclass of simple graphs, we can conclude that Open Packing is NP-complete on $H$-free graphs, a contradiction. This proves the necessary part of Theorem 2.
The sufficiency part of Theorem 2 is held by Theorem 8. ∎
Note that Observation 1, along with Theorem 7, and the fact that Total Dominating Set is NP-complete for (i) Split graphs ($\{2K_{2},C_{4},C_{5}\}$-free graphs) [1], (ii) Bipartite graphs (a subclass of $K_{3}$-free graphs) [16], and (iii) Line graphs (a subclass of $K_{1,3}$-free graphs) [14], imply the following theorem, similar to Theorem 2, for Total Dominating Set.
@@
For $p\geq 4$, let $H$ be a graph on $p$ vertices. Then, Total Dominating Set is polynomial time solvable on the class of $H$-free graphs if and only if $H\in\{pK_{1},(K_{2}\cup(p-2)K_{1}),(P_{3}\cup(p-3)K_{1}),(P_{4}\cup(p-4)K_{1})\}$ unless P = NP.
@@ 3.3 $H$ on three vertices
Note that $3K_{1},(K_{2}\cup K_{1}),P_{3}$, and $K_{3}$ are the only non-isomorphic graph on three vertices. By Theorem 8, an optimal open packing of a $H$-free graph can be found in polynomial time if $H\in\{3K_{1},(K_{2}\cup K_{1}),P_{3}\}$ and Open Packing is NP-complete on $K_{3}$-free graphs by Remark 1. This completes the proof of Theorem 1. ∎
Note that Total Dominating Set is NP-complete for $K_{3}$-free graphs [16]. So, by Theorem 7, a result similar to Theorem 1 for Total Dominating Setholds. This note completes our study on the complexity (P vs NPC) of Total Dominating Set and Open Packing on $H$-free graphs. Next, we study the complexity of these problems on some subclasses of split graphs.
@@ 4.1 $K_{1,r}$-free Split Graphs
We begin this section with a construction to prove that Open Packing is NP-complete on $K_{1,4}$-free split graphs, which is a reduction from Independent Set, a well-known NP-complete problem [12].
@@
Step 1: Subdivide each edge $e=uu^{\prime}$ in $G$ into a three vertex path $ueu^{\prime}$ in $G^{\prime}$. Step 2: Introduce three new vertices, $x$, $y$, and $z$, and two edges, $xy$ and $xz$, in $G^{\prime}$. Step 3: Make $E(G)\cup\{y,z\}$ a clique in $G^{\prime}$.
@@
On the contrary, assume that $\{w,w_{1},w_{2},w_{3},w_{4}\}$ induces a $K_{1,4}$ in $G^{\prime}$ with $w$ as its centre. Then, $\{w_{1},w_{2},w_{3},w_{4}\}$ is an independent set in $G^{\prime}$. So $w\notin I$, since $N_{G^{\prime}}(v)\subseteq C$ is a clique for every vertex $v\in I$, whereas $N_{G^{\prime}}(w)$ contains an independent set on four vertices. This implies that $w\in C$. By Construction 2, for every vertex $v\in C$, $|N_{G^{\prime}}(v)\cap I|\leq 2$, and so $w$ has at most two neighbours in $I$. Therefore, |$|\{w_{1},w_{2},w_{3},w_{4}\}\cap C|\geq 2$, a contradiction to $C$ being a clique and $\{w_{1},w_{2},w_{3},w_{4}\}$ being an independent set in $G^{\prime}$. Thus, $G^{\prime}$ is a $K_{1,4}$-free graph.
The following claims help us prove the guarantee of Construction 2.
@@
Proof of Claim 7 is similar to the proof of Claim 1, and so it is moved to Appendix C.
@@
Observe that in $G^{\prime}$, the graph induced by $C\cup\{x\}$ consists of a complete graph ($G^{\prime}[C]$) on $m+2$ vertices together with a vertex $x$, which is adjacent to exactly two vertices ($y\text{ and }z$) in $C$. So, any two distinct vertices in $C\cup\{x\}$ have a common neighbour. Thus, the claim holds.
@@
The following theorem is held by Construction 2 and the fact that Independent Set is NP-complete on simple graphs [12].
@@
Similar to that of Theorems 5 and 6, it can be proved that on $K_{1,4}$-free split graphs (i) Open Packing parameterized by solution size is W[1]-complete and (ii) Max-Open Packing is hard to approximate within a factor of $N^{(\frac{1}{2}-\epsilon)}$ for any $\epsilon>0$ unless P = NP, where $N$ denotes the number of vertices in a $K_{1,4}$-free split graph (proof of (ii) is given as Theorem 16 in Appendix C).
Next, to complete the dichotomy result of Open Packing in $K_{1,r}$-free split graphs, we prove that a maximum open packing in $K_{1,3}$-free split graphs can be found in polynomial time, and the following claim and lemma are useful in proving it.
@@
Let $S$ be an open packing in a connected split graph $G(C\cup I,E)$ with $S\cap C\neq\emptyset$. Let $u\in S\cap C$. Then, we show that $S\subseteq N[u]$. On the contrary, assume that there exists a vertex $v\in S\setminus N[u]$. Since $C\subseteq N[u]$ and $v\notin N[u]$, $v\notin C$, i.e., $v\in I$. Since $G$ is connected, $deg(v)\geq 1$, and so there exists $w\in C$ such that $vw\in E(G)$. Since $v\notin N(u)$, $w\neq u$. Further, since $w,u\in C$, $w\in N(u)$. This implies that $w\in N(u)\cap N(v)$, a contradiction to $S$ is an open packing in $G$ and $u,v\in S$. Therefore, $S\subseteq N[u]$. Since $S$ is an open packing in $G$, $|S\cap N(u)|\leq 1$. So, $|S|\leq|S\cap N(u)|+1\leq 2$.
@@
Let $I=\{x_{1},x_{2},\ldots,x_{k}\}$. Let $V_{i}=N(x_{i})\cap C$ for $i=1,2,\ldots,k$ and $V_{0}=C\setminus(\cup_{i=1}^{k}V_{i})$. Then, $V_{i}\cap V_{j}=\emptyset$ for every $i,j\in\{1,2,\ldots,k\}$ and $i\neq j$. If not, let $u^{\prime}\in V_{i}\cap V_{j}$ for some $i,j\in\{1,2,\ldots,k\}$ with $i\neq j$. Then, $x_{i},x_{j}\in N(u^{\prime})\cap I$, a contradiction to $|N(u)\cap I|\leq 1$ for every $u\in C$. We prove that in this case $\rho^{o}(G)\in\{2,|I|\}$, and the proof is divided into two subcases.
@@
Note that the assumption in Case 1.2 is the negation of Case 1.1. As we have already proved that $N(x_{i})\cap N(x_{j})=V_{i}\cap V_{j}=\emptyset$ for every $x_{i},x_{j}\in I$, $I$ is an open packing in $G$, and hence $\rho^{o}(G)\geq|I|=k$. Next, we show that $\rho^{o}(G)\leq|I|=k$. On the contrary, assume that $\rho^{o}(G)>k$. Let $S$ be an open packing in $G$ of size $\rho^{o}(G)$. Then, $|S|>|I|=k$, and so $S\cap C\neq\emptyset$. Therefore, by Claim 9, $|S|\leq 2$. This implies that $2\geq|S|>|I|\geq 1$. So, $|S|=2$ and $|I|=k=1$. Then, $V(G)=C\cup\{x_{1}\}$, and by the assumption in Case 1.2, $deg(x_{1})\geq 2$. Let $S=\{a,b\}$. Since $G$ is constituted of a complete graph ($G[C]$) together with a vertex $x_{1}$, which is adjacent to at least two vertices in $C$, for any choice of $a,b\in C\cup\{x_{1}\}$, $a$ and $b$ will have a common neighbour in $G$, a contradiction to $S$ is an open packing of size two in $G$. Therefore, for every open packing $S$ in $G$, $|S|\leq|I|=k$. This implies that $\rho^{o}(G)\leq|I|$. Since we have proved that $I$ is an open packing in $G$, $\rho^{o}(G)=|I|$.
@@
If there exist $x,y\in I$ such that $N(x)\cap N(y)=\emptyset$, then $\{x,y\}$ is an open packing of size two in $G$, and so $\rho^{o}(G)=2$. Also, if there exists a vertex $x\in I$ such that $deg(x)=1$, then it is evident that $n[x]$ is an open packing of size two. Thus, $\rho^{o}(G)=2$.
@@
Note that the assumption in Case 2.2 is the negation of Case 2.1. We prove that $\rho^{o}(G)=1$ in this case. As we have already proved that $\rho^{o}(G)\leq 2$, it is enough to rule out the case $\rho^{o}(G)=2$. Since $V(G)\subseteq C\cup I$, where (i) any two vertices in $I$ have a common neighbour, (ii) every vertex $x\in I$ has at least two neighbours in $C$, and (iii) $G[C]$ is a complete graph, it is evident that any two vertices in $V(G)$ have a common neighbour in $G$, and so $\rho^{o}(G)<2$. Therefore, $\rho^{o}(G)=1$.
Hence, we have shown every possible value of $\rho^{o}(G)$. Note that every $K_{1,3}$-free split graph $G$ falls under Case 1 or 2. It can be easily verified in polynomial time whether a $K_{1,3}$-free split graph $G$ falls under Case 1 or Case 2. Also, the conditions given in Cases 1.1, 1.2, 2.1, and 2.2 are polynomial time verifiable. Hence, the open packing number of a $K_{1,3}$-free split graph can be found in polynomial time. ∎
@@ 4.2 $I_{r}$-split graphs
We observed that a minor modification in Corneil and Perl’s reduction [1] would show that Total Dominating Set is NP-complete on $I_{2}$-split graphs. We extended this by proving a dichotomy that Total Dominating Set in $I_{r}$-split graphs is (i) NP-complete for $r\geq 2$ and (ii) polynomial time solvable for $r=1$. In addition, we prove that Open Packing is (i) NP-complete on $I_{r}$-split graphs for $r\geq 3$ and (ii) polynomial time solvable on $I_{r}$-split graphs for $r\leq 2$. We begin the section with Construction 3 to prove that Total Dominating Set is NP-complete on $I_{r}$-split graphs for $r\geq 2$ through a reduction from $r$-Hitting Set problem, which is known to be NP-complete for $r\geq 2$ [12].
@@
Figure 8: (a) An instance $(U,\mathcal{W})$ of 3-Hitting Set problem and (b) a $I_{3}$-split graph $G(C\cup I,E)$ produced by Construction 3, where $C=U\cup\{x_{1},x_{2},x_{3}\}$ is a clique in $G$ and $I=\{z_{W_{1}},z_{W_{2}},z_{W_{3}},z_{W_{4}},y\}$ is an independent set in $G$ with $deg_{G}(z_{W_{i}})=deg_{G}(y)=3$ for every $i=1,2,3,4$ and $V(G)=C\cup I$.
An example of Construction 3 for the case $r=3$ is given in Fig. 8. Note that $V(G)=C\cup I$, where $C=U\cup\{x_{1},x_{2},\ldots,x_{r}\}$ and $I=\{z_{W}\,:\,W\in\mathcal{W}\}\cup\{y\}$ and $E(G)=\{uv:u,v\in U\cup\{x_{1},x_{2},\ldots,x_{r}\}\}\cup\{uz_{W}\,:\,u\in U,W\in\mathcal{W}\text{ and }u\in W\}\cup\{x_{i}y\,:\,1\leq i\leq r\}$. So, $|V(G)|=|U|+|W|+r+1$ and $|E(G)|=\binom{|U|+r}{2}+(r\cdot|W|)+r$. Hence, the graph $G$ can be constructed in polynomial time with respect to $|U|,|\mathcal{W}|$, and $r$. By the construction of $G$, it is clear that $V(G)=C\cup I$ is a clique-independent set partition of the vertex set of $G$. So, $G$ is a split graph. Since the size of every set $W\in\mathcal{W}$ is $r$, $deg_{G}(z_{W})=r$ by the construction of $G$. Further, since $N_{G}(y)=\{x_{1},x_{2},\ldots,x_{r}\}$, $deg_{G}(y)=r$. This implies that $deg_{G}(u)=r$ for every $u\in I$. Thus, $G(C\cup I,E)$ is an $I_{r}$-split graph.
@@
Let $D$ be a total dominating set of size $k+1$ in $G$. Then, $D\cap\{x_{1},x_{2},\ldots,x_{r}\}\neq\emptyset$ because $N_{G}(y)=\{x_{1},x_{2},\ldots,x_{r}\}$ and $D\cap N_{G}(y)\neq\emptyset$. Let $D^{\prime}=D\cap U$. Since $D\cap\{x_{1},x_{2},\ldots,x_{r}\}\neq\emptyset$ and $U\cap\{x_{1},x_{2},\ldots,x_{r}\}=\emptyset$, $|D^{\prime}|=|D\cap U|\leq k$. Next, we show that $D^{\prime}$ is a hitting set of size at most $k$ of $(U,\mathcal{W})$. Let $W\in\mathcal{W}$. Then, since $D$ is a total dominating set, there exists a vertex $u\in D$ such that $uz_{W}\in E(G)$. Also, since $N_{G}(z_{W})\subseteq U$ by the construction of $G$, $u\in D\cap U=D^{\prime}$. So, by the construction of $G$, $u\in W$. Since $W$ is an arbitrary set in $\mathcal{W}$, for every $W\in\mathcal{W}$, there exists an element $u\in D^{\prime}$ such that $u\in W$. Hence, $D^{\prime}$ is a hitting set of size at most $k$ in $(U,\mathcal{W})$. ∎
@@
The following theorem completes the dichotomy result for Total Dominating Set on $I_{r}$-split graphs.
@@
Let $N(I)=\{u\}$. Since $I$ is an independent set in $G$, $u\in C$. Since $C$ is a clique and $N(I)=\{u\}$, every vertex in $V(G)\setminus\{u\}=(C\cup I)\setminus\{u\}$ is adjacent to $u$. Let $v\in I$. Then, since $N(I)=\{u\}$, $vu\in E(G)$. This implies that every vertex in $V(G)=(V(G)\setminus\{u\})\cup\{u\}$ is adjacent to a vertex in $\{u,v\}$. So, $\{u,v\}$ is a total dominating set in $G$, and hence $\gamma_{t}(G)\leq 2$. Since $\gamma_{t}(H)\geq 2$ for every graph $H$ that admits a total dominating set, $\gamma_{t}(G)=2$.
@@
Since $I$ is an independent set in $G$, $N(I)\subseteq C$. Let $u,u^{\prime}\in N(I)$. Then, since $C$ is a clique and $u,u^{\prime}\in N(I)\subseteq C$, every vertex in $C$ is adjacent to $u$ or $u^{\prime}$. Note that by the definition of $N(I)$, every vertex in $I$ is adjacent to some vertex in $N(I)$. Thus, every vertex in $V(G)=C\cup I$ is adjacent to a vertex in $N(I)$, and so $N(I)$ is a total dominating set. This implies that $\gamma_{t}(G)\leq|N(I)|$. Recall that $\gamma_{t}(G)\geq|N(I)|$. Hence, $\gamma_{t}(G)=|N(I)|$.
Thus, by Cases 1 and 2, $\gamma_{t}(G)=\max\{2,|N(I)|\}$. Note that $|N(I)|$ can be found in linear time. So, an optimal total dominating set of $G$ can be found in linear time. ∎
Next, we prove the dichotomy result on Open Packing in $I_{r}$-split graphs. Firstly, we show that Open Packing is NP-complete on $I_{r}$-split graphs for $r\geq 3$ through a reduction from $r$-Dimensional Matching problem, which is known to be NP-complete for $r\geq 3$ [12].
@@
Input: A collection of sets $X_{1},X_{2},\ldots,X_{r}$ such that $|X_{i}|=q$ for $i=1,2,\ldots,r$, for some $q\in\mathbb{N}$, and a non-empty set $M\subseteq\prod\limits_{i=1}^{r}X_{i}$.